ambreerheinberger:

Developer wants to enclose a rectangular grassy lot that borders a city street for parking. If the developer has 308 feet of fencing is not fence the side along the street what is the largest area that can be enclosed

justjm:

Here is what we know: •Perimeter is 308 ft •Fencing is rectangular in figure but only encloses 3 sides. Let x=width and y=length  We can therefore form two systems of equations: \((1)...308=2x+y\) \((2)...A=xy\) Now we are supposed to find the largest area enclosed using these requirements. You must first solve for y in equation 1 so that you can use equation 2 to find the max area. \((1)...308=2x+y\) \(308-2x=y\) ∴\(A=x(308-2x)\) \(A=308x-2x^2\) Now, to find the maximum, you must first find the critical points, where \(\frac{dA}{dx}=0 ~\text{and}~DNE \). Take the derivative. \(A=308x-2x^2\) \(\frac{dA}{dx}=308-4x\) \(0=308-4x\) \(x=\frac{-308}{-4}=77\) \(\frac{1}{0}=308-4x\) There is no such value of x that yields an indeterminate value. Hence, your critical point is x=77. Now, you must test whether this is a maximum, minimum, or neither. Let's use the first derivative test. \(•\frac{dA}{dt}|_{t<77}\) \(\frac{dA}{dx}|_{t=70}=28\) \(•\frac{dA}{dx}|_{t>77}\) \(\frac{dA}{dt}|_{t=80}=-12\) ∴Since \(\frac{dA}{dt}\) changes from positive to negative at x=77, when the width is 77, the area reaches a maximum. But that is not the full question. We are asked to find the AREA. So since we know x=77, find y. \(308=2(77)+y\) \(y=154\) \(A_{max}=(77)(154)=11,858~ft^2\)

justjm:

Or you could just take that area function and find the vertex, it doesn't matter